### v1

@@ Abstract
Method: This SLR was conducted according to the Kitchenham and Charters’ guidelines. We used both manual and automatic search strategies without putting any time limitation for searching the relevant papers.
Results: We selected 63 primary studies for data extraction, analysis, and synthesis based on our pre-defined inclusion and exclusion criteria. From the results of the data analysis, we classified the extracted data into 4 categories based on the data extraction source, categorized the proposed recommendation systems to fit into a taxonomy, and identified the key advantages and limitations of these systems. Our results revealed that human factors play a major role in CSE task recommendation. Further, we identified five popular task types recommended, popular platforms, and their features used in task recommendation. We also provided recommendations for future research directions.
Conclusion: This SLR provides insights into current trends, gaps, and future research directions in CSE task recommendation systems.
@@ 1 Introduction
Crowdsourcing is a distributed problem-solving approach that combines human and machine computation. It involves organizations outsourcing work to an undefined networked labor pool through open calls for participation [1]. Crowdsourced Software Engineering (CSE) is derived from the crowdsourcing concept. CSE utilizes open calls to recruit global online labor in various software engineering tasks, including requirements extraction, design, coding, and testing [1]. It reduces time-to-market by enabling parallelism, lowering costs and defect rates, and enabling flexible development capabilities [2][3][4]. Mao et al. state that there is a dramatic rise in recent work on the use of crowdsourcing in software engineering based on their conducted pilot study [1].
@@
Popular crowdsourcing platform, Topcoder (www.topcoder.com), has hosted over 136,202 challenges in software design, development, and data science, with total prizes exceeding US\$191,873,091 [5]. uTest (www.utest.com) has a freelancer pool of more than 1 million testers worldwide who test new apps for device compatibility, conduct functionality testing, and perform usability inspections [6]. As of April 2024, StackOverflow (stackoverflow.com) has over 24 million programming related questions and 36 million answers [7]. Major companies like Netflix, Microsoft, Facebook, and Google frequently offer bug bounties, particularly for security vulnerabilities [8]. With such large community involvement, it is crucial to recommend the most suitable tasks to the contributors.
@@
Developers often encounter various technical and social barriers when attempting to onboard crowdsourcing projects [9]. Unsuccessful onboarding not only affects developers but also impacts the projects themselves. It increases the time and effort new developers spend searching for suitable projects and learning [10]. Further, online crowdsourced software development operates as a globally distributed collaborative effort. It involves a diverse range of contributors with varying personalities, educational backgrounds, and levels of expertise. These contributors work from different locations across multiple time zones and undertake various development tasks, such as bug fixing, code testing, or documentation, driven by their individual motivations. Hence, these contributors may struggle to find suitable projects or tasks among the thousands available in these communities [11]. Various recommendation systems have been introduced in the literature to provide personalized task recommendations to software practitioners on crowdsourcing platforms. Providing such recommendations results in increased worker satisfaction, lower churning rate after onboarding, and increased task quality.
This SLR provides an in-depth understanding of existing literature on providing personalised task recommendations to software practitioners in crowdsourcing platforms. The key contributions of this SLR are as follows:
@@
  • Identifying, analyzing, extracting, and synthesizing 63 highly relevant primary studies.
@@
  • Identifying key limitations in the existing studies
@@
The rest of the paper is structured as follows. Section 2 provides a summary of the existing research. Section 3 describes the method used for this systematic literature review. Section 4 presents the demographic information of the presented studies and answers the research questions of this SLR. The threats to validity are discussed in Section 5. The insights into future research directions are discussed in Section 6. Finally, the conclusions of this study are presented in Section 7.
2 Existing Literature Reviews
Zhen et al. conducted an SLR that focuses on the software tasks in the crowdsourcing domain at a high level, along with the tasks in non-software engineering domains such as data entry, writing, editing, etc. [12]. Our SLR provides a comprehensive, in-depth analysis of personalized task recommendations to software practitioners in crowdsourcing platforms, focusing on recommending a task to a particular software practitioner in a crowdsourcing platform. Further, the existing SLR is conducted on the studies published from 2010 to 2019, whereas our SLR has no time duration restriction. Our search string significantly differs from the one used in [12]. They identified 120 relevant studies and selected 70 most relevant studies for their SLR. Only one primary study overlaps with the selected primary studies of our SLR. Regarding the research questions, our RQ1, RQ2, and RQ4 partially overlap with the existing literature review’s RQ3, RQ2, and RQ4, respectively. Moreover, we focus on assigning a suitable task to a particular software practitioner, whereas the majority of primary studies in the existing SLR focus on assigning a practitioner to a task. Since the existing SLR is a high-level study on the software engineering domain, the results include those unrelated to the software engineering field. Hence, there is a significant difference between our SLR and the existing SLR.
Mao et al. performed a survey that focuses on how crowdsourcing is used in different stages of the software development life cycle. Further, it provides definitions for crowdsourced software engineering, crowdsourcing practices in software engineering, commercial platforms, and relevant case studies. The survey is not based on any research questions as it is a broad analysis. They conducted the survey on the studies published from January 2006 to March 2015, whereas our SLR has no time duration restriction. They surveyed 210 publications out of 476 studies identified during the initial stage. There are no overlapping primary studies with our selected primary studies set for this SLR. Our search string significantly differs from the search string used in this survey. RQ4 in our SLR overlaps with the section “4.1 commercial platforms” in the survey. However, our SLR is focused on recommending crowdsourced software engineering tasks to software practitioners, whereas this survey is a broad analysis of crowdsourced software engineering. Hence, there is a significant difference in this survey with our SLR.
@@ 3.1 Research questions
The research questions (RQs) were developed by the PICOC approach introduced by Petticrew and Roberts [14], explained in Kitchenham and Charters’ guidelines [13]. The PICOC (Population, Intervention, Comparison, Outcomes, Context) table for this SLR is available in Table 1. The intervention is the software technologies that address the issue, the population is the people affected by the intervention, the comparison is the technology used in software engineering that is being compared to the intervention, the outcomes are the factors of importance to practitioners, context refers to the context in which the comparison takes place, the people participating in the research, and the tasks being carried out.
@@
Population Software practitioners in crowdsourcing platforms (eg: developers, testers)
@@
RQ 1: How has the task recommendation for software repositories been done so far and what is its current status?
This RQ aims to investigate the methods that have been introduced to recommend software tasks in crowdsourcing platforms. Further, we aim to categorize the identified methods introduced in the selected primary studies. We also aim to examine the advantages and limitations of the selected primary studies to specify the current strengths and limitations of the area.
@@
RQ2: What are the common types of tasks that can be recommended to software practitioners?
This RQ focuses on the different types of tasks that can be recommended to software practitioners who are in crowdsourcing platforms. This research question provides insights into the areas of current focus within the software crowdsourcing industry.
RQ 3: To what extent are human factors used to recommend tasks in software repositories?
@@
According to Dutra et al., human factors, also known as soft factors, human aspects, social factors, or non-technical factors, refer to the physical or cognitive features, or social behaviors of individuals [15]. Human factors cover a wide range of aspects including personality, motivation, cognition, communication, decision-making, and cohesion among developers as they carry out their tasks [16]. Avison et al. [17] state that “Failure to include human factors may explain some of the dissatisfaction with conventional information systems development methodologies; they do not address real organizations". Since software development is a human-centered process, human factors significantly impact both the process and its performance [18]. Hence this RQ examines the human factors that are currently being considered in the software crowdsourcing task recommendation systems. It also identifies the human factors suggested to be included as future improvements in the selected primary studies.
RQ4: What software repositories are currently being used for task recommendations?
@@
This RQ identifies the different software crowdsourcing platforms that are currently being used for task recommendations. It further analyses the scale of the projects that are used to recommend the tasks. The answers to this research question provide insights into popular crowdsourcing platforms.
@@
RQ5: Which features of software repositories can be used to recommend tasks?
@@
This RQ identifies the features of the crowdsourcing platforms that can be used to recommend tasks. It also identifies the data items that are currently being extracted from those crowdsourcing platforms. The answers to this research question provide a comprehensive set of features and data attributes that can be used in future studies.
@@ 3.2.1 Search method
The automatic and manual searches were carried out to find the relevant primary studies for this SLR. Scopus, IEEE Xplore, ACM Digital Library, Wiley, Science Direct, and Springer Link scientific databases were queried during the automatic search using the search query defined in Section 3.2.2. Then, a manual search was done via backward and forward snowballing on the finalized relevant primary studies retrieved by the automatic search.
@@ 3.2.2 Search terms
The search query was formulated according to the Kitchenham and Charters’ guidelines [13]. The PICOC approach (see Table 1) was used to select the key search terms. The search query was finalized after a series of pilot searches and examining the resulting studies. In some studies, it was observed that generic terms like “recommendation algorithm” and “recommendation system” were used instead of any crowdsourcing-related terms within the title, abstract, and keywords, e.g., “good first issue recommendation system”. Hence, the aforementioned two terms were also added with the alternative search terms for “software crowdsourcing”. The final search query was formulated by connecting the alternative search terms with “OR” and “AND”, which are mentioned below. The alternative search terms used in the query are mentioned in Table 2. The wild cards were used except in the Science Direct database to cover a large number of alternative search terms. In Science Direct’s advanced search, only 8 operators (e.g., AND/OR) are allowed per query. Hence, the combinations of the search terms were used to maintain consistency among the search queries used in different databases. Further, the search was not restricted by any particular period.
@@ 3.3 Inclusion and exclusion criteria
I02 Studies that are relevant to recommending software engineering-related tasks to a software practitioner in a crowdsourcing platform
@@
E02 Papers that are not written in English language
E03 Short papers which have less than 4 pages
@@ 3.4 Study Selection
Phase 2: During the second phase, the potential studies set was screened by reading the title, abstract, and keywords. The papers which we had doubts regarding the relevancy were transferred to the next round of screening. The final count of potential studies at the end of the first screening round was 132.
Phase 3: During the third phase, the filtering was done by reading the full-text papers of the potentially relevant studies found during the second phase. The data extraction of the primary studies was also done during this stage. After this filtering round, 51 papers were left as the primary studies.
Phase 4: The snowballing was done during the last phase for the finalized primary studies. Both backward and forward snowballing processes were used, and 9 and 12 potential studies were found, respectively. Then, the first filtration for those studies was done based on the title, abstract, and keywords. Then, 6 papers were left for backward snowballing, and 8 were left for forward snowballing at the end of filtration round 1. Then during the next filtration round, we read the full-text papers of those studies. After the second filtration, the total number of relevant primary studies finalized by backward snowballing was 5, and forward snowballing was 7.
@@
Database Initial Results Count Filtration 1 Filtration 2
Automatic Search Scopus 836 85 5
IEEE Xplore 98 26 19
ACM Digital Library 32 9 15
@@
Total - Automatic Search 1368 132 51
@@
Final Paper Count 63
@@ 3.5 Data extraction and synthesis
The data items extracted from the selected primary studies to answer the research questions of the SLR are presented in Table 7. The data items were retrieved and categorized based on each research question. The extracted data were collected in a Microsoft Excel Spreadsheet file. Data synthesis of the extracted data was done by thematic analysis [20]. This approach is often employed in systematic literature reviews (SLRs). It can be defined as an analysis of an extensive amount of structured, extracted data from individual studies to integrate and synthesise their main findings.

### v2

@@ Abstract
Method: This SLR was conducted according to the Kitchenham and Charters’ guidelines. We used manual and automatic search strategies without putting any time limitation for searching the relevant papers.
Results: We selected 65 primary studies for data extraction, analysis, and synthesis based on our predefined inclusion and exclusion criteria. Based on our data analysis results, we classified the extracted information into four categories according to the data acquisition sources: Software Practitioner’s Profile, Task or Project, Previous Contributions, and Direct Data Collection. We also organized the proposed recommendation systems into a taxonomy and identified key advantages, such as increased performance, accuracy, and optimized solutions. In addition, we identified the limitations of these systems, such as inadequate or biased recommendations and lack of generalizability. Our results revealed that human factors play a major role in CSE task recommendation. Further, we identified five popular task types recommended, popular platforms, and their features used in task recommendation. We also provided recommendations for future research directions.
Conclusion: This SLR provides insights into current trends, gaps, and future research directions in CSE task recommendation systems such as the need for comprehensive evaluation, standardized evaluation metrics, and benchmarking in future studies, transferring knowledge from other platforms to address cold start problem.
@@ 1 Introduction
Crowdsourcing is a distributed problem-solving approach that combines human and machine computation. It involves organizations outsourcing work to an undefined networked labour pool through open calls for participation [1]. Crowdsourced Software Engineering (CSE) is derived from the crowdsourcing concept. CSE utilizes open calls to recruit global online labour in various software engineering tasks, including requirements extraction, design, coding, and testing [1]. It reduces time-to-market by enabling parallelism, lowering costs and defect rates, and enabling flexible development capabilities [2][3][4]. Mao et al. state that there is a dramatic rise in recent work on the use of crowdsourcing in software engineering based on their conducted pilot study [1].
@@
Popular crowdsourcing platform Topcoder 111 https://www.topcoder.com has hosted over 136,202 challenges in software design, development, and data science, with total prizes exceeding US\$191,873,091 [5]. uTest 222 https://www.utest.com has a freelancer pool of more than 1 million testers worldwide who test new apps for device compatibility, conduct functionality testing, and perform usability inspections [6]. As of April 2024, StackOverflow 333 stackoverflow.com has over 24 million programming-related questions and 36 million answers [7]. Major companies like Netflix, Microsoft, Facebook, and Google frequently offer bug bounties, particularly for security vulnerabilities [8]. With such extensive community involvement, it is crucial to recommend the most suitable tasks to the contributors.
@@
Developers often encounter various technical and social barriers when attempting to onboard crowdsourcing projects [9]. Unsuccessful onboarding not only affects developers but also impacts the projects themselves. It increases the time and effort new developers spend searching for suitable projects and learning [10]. Further, online crowdsourced software development operates as a globally distributed collaborative effort. It involves a diverse range of contributors with varying personalities, educational backgrounds, and levels of expertise. These contributors work from different locations across multiple time zones and undertake various development tasks, such as bug fixing, code testing, or documentation, driven by their individual motivations. Hence, these contributors may struggle to find suitable projects or tasks among the thousands available in these communities [11]. Various recommendation systems have been introduced in the literature to provide personalized task recommendations to software practitioners on crowdsourcing platforms. Providing such recommendations results in increased worker satisfaction, lower churning rate after onboarding, and improved task quality.
We retrieved relevant studies from 6 popular digital databases, including Scopus, IEEE Xplore, and ACM digital library, where we found 6, 20, 15, 5, 3, and 4 primary studies, respectively. After forward and backward snowballing, the total number of primary studies led to 65. We found an existing SLR and a survey that are slightly relevant to the focused domain of this study. However, the SLR by Zhen et al. [12] focuses broadly on software as well as non-software tasks in crowdsourcing, while our SLR provides an in-depth analysis of personalized CSE task recommendations for software practitioners. Similarly, Mao et al. [1]’s survey broadly analyzes crowdsourced software engineering practices without specific research questions or overlapping studies, whereas our SLR focuses specifically on recommending tasks to practitioners.
This SLR aims to consolidate existing information and provide an in-depth understanding of the existing literature on providing personalized task recommendations to software practitioners on crowdsourcing platforms. The findings of this study offer insights for both practitioners and researchers focused on enhancing the CSE task recommendation systems while also laying the groundwork for future research in this field. The key contributions of this SLR are as follows:
@@
  • Identifying, analyzing, extracting, and synthesizing 65 highly relevant primary studies.
@@
  • Identifying key advantages and limitations in the existing studies.
@@
The key results of our study revealed that the data used to recommend tasks were sourced from 4 sources: software practitioner’s profile, task or project itself, previously contributed tasks or projects available in the crowdsourcing platform and direct data collection via a customized platform. The existing software task recommendation methods can be mainly categorized as content-based, collaborative filtering, and hybrid approaches. Content-based approaches are the most popular category. The existing CSE task recommendation systems offer key advantages such as increased performance, accuracy, optimized solutions, and the ability to provide more comprehensive and diverse recommendations. However, they also have three main limitations: limitations in the dataset, approach, and evaluation. Commonly recommended tasks in CSE platforms, such as GitHub, Topcoder, and Utest, include repository contributions, issue fixes, testing tasks, Good First Issues (GFI), and answering questions. Among the various features used in existing systems, issue or task descriptions are the most widely utilized for data extraction. Key future research directions include the need for integrating more human factors into these systems, transferring knowledge from other platforms to overcome data sparsity, and exploring the possibility of integrating these recommendation systems with development workflow tools.
The rest of the paper is structured as follows. Section 2 describes the method used for this systematic literature review. Section 3 presents the demographic information of the presented studies and answers the research questions of this SLR. The threats to validity are discussed in Section 4. Section 5 provides a summary of the existing research. The insights into future research directions are discussed in Section 6. Finally, the conclusions of this study are presented in Section 7.
@@ 2.1 Research questions
The research questions (RQs) were developed by the PICOC approach introduced by Petticrew and Roberts [14], explained in Kitchenham and Charters’ guidelines [13]. This approach was originally developed in the field of Health Sciences but has been adapted for other domains, such as Information Technology. It helps streamline the identification of key elements essential for formulating clear research questions and selecting keywords for search string construction. The PICOC (Population, Intervention, Comparison, Outcomes, Context) table for this SLR is available in Table 1. The intervention is the software technologies that address the issue, the population is the people affected by the intervention, the comparison is the technology used in software engineering that is being compared to the intervention, the outcomes are the factors of importance to practitioners, context refers to the context in which the comparison takes place, the people participating in the research, and the tasks being carried out.
@@
Population Software practitioners in crowdsourcing platforms (e.g.; developers, testers)
@@
RQ 1: How has the task recommendation in crowdsourced software engineering been done so far and what is its current status?
This research question is studied under four sub-questions mentioned below. It aims to investigate the state-of-the-art methods in task recommendation for CSE. It involves exploring the data sources and types of data extracted for task recommendations, categorizing the existing methods, and examining their advantages and limitations to understand the current landscape. It establishes the current state and methodologies of task recommendation systems, setting the context for subsequent RQs. Further, it lays the foundation for identifying opportunities for improvement.
RQ1.1: What data sources are used and what type of data are extracted to recommend tasks?
RQ1.2: What are the existing software task recommendation methods?
RQ1.3: What are the advantages of the existing task recommendation systems?
RQ1.4: What are the limitations of the existing task recommendation systems?
@@
RQ2: What are the common types of tasks that are recommended to software practitioners?
This RQ focuses on understanding the types of tasks commonly recommended to practitioners in CSE platforms. It highlights the areas where task recommendation systems are currently concentrated and provides insights into the scope and diversity of tasks addressed by existing systems, offering a clearer picture of the industry’s focus. This RQ builds on RQ1 by narrowing the focus to the types of tasks addressed by these systems, highlighting industry priorities.
RQ 3: To what extent are human factors used to recommend CSE tasks?
@@
According to Dutra et al., human factors, also known as soft factors, human aspects, social factors, or non-technical factors, refer to the physical or cognitive features, or social behaviors of individuals [15]. Human factors cover a wide range of aspects, including personality, motivation, cognition, communication, decision-making, and cohesion among developers as they carry out their tasks [16]. Avison et al. [17] state that “Failure to include human factors may explain some of the dissatisfaction with conventional information systems development methodologies; they do not address real organizations". Since software development is a human-centered process, human factors significantly impact both the process and its performance [18]. Given the human-centric nature of software development, understanding how human factors are integrated into CSE task recommendation systems helps identify gaps in addressing the social and cognitive needs of developers. Hence, this RQ examines the human factors that are currently being considered in the software crowdsourcing task recommendation systems. It also identifies the human factors suggested to be included as future improvements in the selected primary studies. This RQ explores the integration of human factors, a critical dimension that complements the technical aspects addressed in RQ1 and RQ2.
RQ4:What platforms are currently being used to recommend crowdsourced software engineering tasks?
@@
This RQ identifies the different software crowdsourcing platforms currently being used for task recommendations. It further analyses the scale of the projects used to recommend tasks. The answers to this research question provide insights into popular crowdsourcing platforms and their popularity and suitability for different task recommendation needs.
@@
RQ5:Which features of platforms can be used to recommend crowdsourced software engineering tasks?
@@
This RQ identifies the features of crowdsourcing platforms that can be used to recommend tasks and the data items currently being extracted from them. The answers to this research question provide a comprehensive set of features and data attributes, offering actionable insights for system design and improvement in future studies.
@@ 2.2.1 Search method
The automatic and manual searches were carried out to find the relevant primary studies for this SLR. Scopus, IEEE Xplore, ACM Digital Library, Wiley, Science Direct, and Springer Link scientific databases were queried during the automatic search using the search query defined in Section 2.2.2. Then, backward and forward snowballing was done on the finalized relevant primary studies retrieved by the automatic search.
@@ 2.2.2 Search terms
The search query was formulated according to the Kitchenham and Charters’ guidelines [13]. The PICOC approach (see Table 1) was used to select the key search terms. The search query was finalized after a series of pilot searches and examining the resulting studies. In some studies, it was observed that generic terms like “recommendation algorithm” and “recommendation system” were used instead of any crowdsourcing-related terms within the title, abstract, and keywords, e.g., “good first issue recommendation system”. Hence, the aforementioned two terms were also added with the alternative search terms for “software crowdsourcing”. The final search query was formulated by connecting the alternative search terms with “OR” and “AND”, which are mentioned below. The alternative search terms used in the query are mentioned in Table 2. The wild cards were used except in the Science Direct database to cover a large number of alternative search terms. In Science Direct’s advanced search, only 8 operators (e.g., AND/OR) are allowed per query. Hence, the combinations of the search terms were used to maintain consistency among the search queries used in different databases. Further, the search was not restricted by any particular period. Finally, we randomly selected 8-10 papers from each digital database to confirm the credibility of our search string and validate that the included studies were the most pertinent to our review.
@@ 2.3 Inclusion and exclusion criteria
I02 Studies that are focused on recommending software engineering-related tasks to a software practitioner in a crowdsourcing platform
I03 Study is published in one of the digital libraries we queried for this study
@@
E02 Studies that are focused on recommending a software practitioner to do a software engineering-related task published in a crowdsourcing platform
E03 Papers that are not written in English language
E04 Short papers which have less than 4 pages
E03 Papers that do not propose a recommendation system based solution
@@ 2.4 Study Selection
Phase 2: During the second phase, the potential studies set was screened by reading the title, abstract, and keywords. The papers which we had doubts regarding the relevancy were transferred to the next round of screening. The final count of potential studies at the end of the first screening round was 142.
Phase 3: During the third phase, the filtering was done by reading the full-text papers of the potentially relevant studies found during the second phase. If the proposed solution of the study is not to recommend a software engineering-related task to a software practitioner or instead it is to recommend a software practitioner to do a task in a CSE platform, such studies were excluded in this stage after a careful evaluation of the proposed solution. The data extraction of the primary studies was also done during this stage. After this filtering round, 51 papers were left as the primary studies.
Phase 4: The snowballing was done during the last phase for the finalized primary studies. Both backward and forward snowballing processes were used, and 9 and 12 potential studies were found, respectively. Then, the first filtration for those studies was done based on the title, abstract, and keywords. Then, 6 papers were left for backward snowballing, and 8 were left for forward snowballing at the end of filtration round 1. Then during the next filtration round, we read the full-text papers of those studies. After the second filtration, the total number of relevant primary studies finalized by backward snowballing was 5, and forward snowballing was 7. The finalized set of primary studies are available in Appendix A.
@@
Database Initial Results Count Filter by Title, Abstract, and Keywords Filter by Full text reading
Automatic Search Scopus 836 85 6
IEEE Xplore 98 26 20
ACM Digital Library 32 19 15
@@
Total - Automatic Search 1368 142 53
@@
Final Paper Count 65
@@
2.5 Quality assessment
We implemented a five-point scoring system to evaluate the quality of the selected primary studies based on five predefined questions. The scores range from very poor (1) and inadequate (2) to moderate (3), good (4), and excellent (5). This approach to quality assessment is widely used in SLRs [20, 21]. Our quality assessment (QA) criteria were derived from Kitchenham’s guidelines [13] and are outlined as follows:
  • QA1:Are the aims clearly stated?
  • QA2: Are the measures used clearly defined?
  • QA3: Is the solution clearly defined?
  • QA4: Is there a clear outcome and results analysis reported?
  • QA5: Are study limitations and possible future work adequately described?
The results of this assessment are presented in Appendix B. Our assessment revealed that among the 65 included studies, 14 were of good quality, 39 were of average quality, and 12 were of poor quality. No studies were excluded based on the quality assessment score as recommending CSE tasks to practitioners is an emerging research area, and our selection was already constrained. We did not consider the citation count of each paper as a quality criterion to minimize publication bias and to ensure fairness toward recently published studies.
@@ 2.6 Data extraction and synthesis
The data items extracted from the selected primary studies to answer the research questions of the SLR are presented in Table 7. The data items were retrieved and categorized based on each research question. The extracted data were collected in a Microsoft Excel Spreadsheet file. Data synthesis of the extracted data was done by thematic analysis [22, 23]. This approach is often employed in systematic literature reviews (SLRs). It can be defined as an analysis of an extensive amount of structured, extracted data from individual studies to integrate and synthesise their main findings. The analysis was conducted by the first researcher, while the second researcher reviewed all extracted themes to validate them and identify any additional potential themes. We performed the thematic analysis using the five steps outlined by [23].
  1. 1.
    Familiarizing with data: The extracted data items, such as task types (D08), advantages (D06), limitations (D07), data extracted from the user (D02), and data extracted from the task (D03), were carefully read and examined to develop an initial understanding.
  2. 2.
    Generating initial codes: Lists of advantages, limitations, data extraction sources, etc., were compiled for each solution. In some cases, the original papers were revisited to ensure accuracy (Fig. 3.A).
  3. 3.
    Searching for themes: The related initial codes from the previous step were grouped to form potential themes for each data item (Fig. 3.B).
  4. 4.
    Reviewing themes: Identified themes were compared and analyzed, assessing whether some needed to be merged or removed.
  5. 5.
    Defining and naming themes: Finally, clear and meaningful names were assigned to each identified data extraction source, advantage, limitation, etc., ensuring they accurately represented the findings (Fig. 3.C).
Figure 3 shows the application of thematic analysis to analyze various advantages and identify the "Making more comprehensive diverse recommendations" theme.
Figure 3: The steps of performing thematic analysis